OpenStudy (anonymous):

Differentiate the following function. h (θ) = 9 csc(θ) + 3 e^θ cot(θ)

OpenStudy (anonymous):

h'(\theta) = -9csc(\theta)cot(\theta) + 3 [e^(\theta)cot(\theta) + e^(\theta)[-csc^2(\theta)] = -9csc(\theta)cot(\theta) + 3e^(\theta)[cot(\theta) - csc^2(\theta)]

OpenStudy (anonymous):

the derivative of e^x is x'e^x -> e^x, and for the trig functions, just memorize their derivatives.

OpenStudy (anonymous):

csc'=-csccot cot'=-csc^2

OpenStudy (anonymous):

ty!!

OpenStudy (anonymous):

\[\int\limits\limits\limits \left(3 e^{\theta } \text{Cot}[\theta ]-9 \text{Cot}[\theta ] \text{Csc}[\theta ]-3 e^{\theta } \text{Csc}[\theta ]^2\right) \, d\theta =\] \[3 e^{\theta } \text{Cot}[\theta ]+9 \text{Csc}[\theta ] \]